# Uniqueness of the inverse source problem for fractional diffusion-wave equations

### v1

@@ 3.4 Proof of the theorems
If $f$ satisfies the regularity condition as described in Yamamoto’s work [Theorem 1, [22]], our methodology for proving Theorem 2.2, based on Proposition 3.6, aligns closely with the approach in [22]. However, in scenarios where $f$ is characterized by a weaker regularity condition, specifically $f\in\mathcal{D}(A)$, a modification is necessary: in equation (3.3) of Yamamoto’s demonstration, $\tilde{v}=Av$ should be substituted with $v^{(k)}=A^{k}v$. Despite these similarities, we opt to detail our proof to illustrate its connection to the subsequent theorems.
@@
$\left|\sum_{n=1}^{\infty}a_{n}\psi_{n}(t)\right|\leq||u(\cdot,t)||_{L^{2}(\gamma)}||v||_{L^{2}(\gamma)}=O(t^{-\infty})\text{ as }t\rightarrow\infty.$
@@
Therefore,
$\sum_{n=1}^{\infty}|a_{n}\lambda_{n}^{-1}\log(1+\lambda_{n})|\leq C||g_{2}||_{L^{2}(\Omega)}\sum_{n=1}^{\infty}\lambda_{n}^{-1/4+\epsilon/2-\sigma_{2}}\log(1+\lambda_{n}).$ (32)
When $n$ is large enough, we have $\lambda_{n}^{-1/4+\epsilon/2-\sigma_{2}}\log(1+\lambda_{n})\leq\frac{c_{\lambda}}{2}n^{-2/d(\sigma_{2}+1/4-\epsilon)}$. By definition of $\sigma_{2}$, we can choose $\epsilon$ such that $20<\epsilon<\sigma_{2}+1/4-d/$ 2. Now the series in equation (32) converges, and we can argue similarly to part (a) to finish the proof. $\square$
@@
Setting $g_{3}=A^{\sigma_{3}}f\in L^{2}(\Omega)$, we get $|a_{n}|\leq C\lambda_{n}^{d/4+\epsilon/2-\sigma_{3}}||P_{n}g_{3}||_{L^{2}(\Omega)}$. It follows that
$\sum_{n=1}^{\infty}|a_{n}\lambda_{n}^{-1}\log(1+\lambda_{n})|\leq C||g_{3}||_{L^{2}(\Omega)}\sum_{n=1}^{\infty}\lambda_{n}^{-1+d/4+\epsilon/2-\sigma_{3}}\log(1+\lambda_{n}).$ (33)
@@
Set $g_{4}=A^{\sigma_{4}}f\in L^{2}(\Omega)$, then
$\sum_{n=1}^{\infty}|a_{n}\lambda_{n}^{-1}\log(1+\lambda_{n})|\leq C||g_{4}||_{L^{2}(\Omega)}\sum_{n=1}^{\infty}\lambda_{n}^{(d-1)/4+\epsilon/2-\sigma_{4}}\log(1+\lambda_{n}).$ (34)

### v2

@@ 3.4 Proof of the theorems
If $f$ satisfies the regularity condition as described in Yamamoto’s work [Theorem 1, [22]], our methodology for proving Theorem 2.2, based on Proposition 3.6, aligns closely with the approach in [22]. However, in scenarios where $f$ is characterized by a weaker regularity condition, specifically $f\in L^{2}(\Omega)$, a modification is necessary: in equation (3.3) of Yamamoto’s demonstration, $\tilde{v}=Av$ should be substituted with $v^{(k)}=A^{k}v$. Despite these similarities, we opt to detail our proof to illustrate its connection to the subsequent theorems.
@@
$\left|\sum_{n=1}^{\infty}a_{n}\psi_{n}(t)\right|\leq||\partial_{\nu}u(\cdot,t)||_{L^{2}(\gamma)}||v||_{L^{2}(\gamma)}=O(t^{-\infty})\text{ as }t\rightarrow\infty.$
@@
Using Cauchy-Schwartz inequality, we obtain
$\sum_{n=1}^{\infty}|a_{n}\lambda_{n}^{-1}\log(1+\lambda_{n})|$ $\leq C\left(\sum_{n=1}^{\infty}||P_{n}g_{2}||_{L^{2}(\Omega)}^{2}\right)^{1/2}\left(\sum_{n=1}^{\infty}\lambda_{n}^{-1/2+\epsilon-2\sigma_{2}}(\log(1+\lambda_{n}))^{2}\right)^{1/2}$ (34)
$=C||g_{2}||_{L^{2}(\Omega)}\left(\sum_{n=1}^{\infty}\lambda_{n}^{-1/2+\epsilon-2\sigma_{2}}(\log(1+\lambda_{n}))^{2}\right)^{1/2}.$
When $n$ is large enough, we have $\lambda_{n}^{-1/2+\epsilon-2\sigma_{2}}(\log(1+\lambda_{n}))^{2}\leq\frac{c_{\lambda}}{2}n^{-2/d(2\sigma_{2}+1/2-2\epsilon)}$. By definition of $\sigma_{2}$, we can choose $\epsilon$ such that $40<\epsilon<\sigma_{2}+1/4-d/$ 4. Now the series in equation (34) converges, and we can argue similarly to part (a) to finish the proof. $\square$
@@
Setting $g_{3}=A^{\sigma_{3}}f\in L^{2}(\Omega)$, we get $|a_{n}|\leq C\lambda_{n}^{d/4+\epsilon/2-\sigma_{3}}||P_{n}g_{3}||_{L^{2}(\Omega)}$. It follows from Cauchy-Schwartz inequality that
$\sum_{n=1}^{\infty}|a_{n}\lambda_{n}^{-1}\log(1+\lambda_{n})|$ $\leq C\left(\sum_{n=1}^{\infty}||P_{n}g_{3}||_{L^{2}(\Omega)}^{2}\right)^{1/2}\left(\sum_{n=1}^{\infty}\lambda_{n}^{-2+d/2+\epsilon-2\sigma_{3}}(\log(1+\lambda_{n}))^{2}\right)^{1/2}$ (35)
$=C||g_{3}||_{L^{2}(\Omega)}\left(\sum_{n=1}^{\infty}\lambda_{n}^{-2+d/2+\epsilon-2\sigma_{3}}(\log(1+\lambda_{n}))^{2}\right)^{1/2}.$
@@
Setting $g_{4}=A^{\sigma_{4}}f\in L^{2}(\Omega)$, then similar to equation (35) we obtain
$\sum_{n=1}^{\infty}|a_{n}\lambda_{n}^{-1}\log(1+\lambda_{n})|\leq C||g_{4}||_{L^{2}(\Omega)}\left(\sum_{n=1}^{\infty}\lambda_{n}^{(d-1)/2+\epsilon-2\sigma_{4}}(\log(1+\lambda_{n}))^{2}\right)^{1/2}.$ (36)